OpenStudy (anonymous):

how do you find the solutions for (3x-1)^2=(x-1)^2

OpenStudy (ajprincess):

Take square root both sides. This will give u 3x-1=x-1 Nw try to solve for x:) Does that help?

OpenStudy (anonymous):

yeah i know that but I don't get how to come to the answers x=0 or x=0,5

OpenStudy (ajprincess):

\((3x-1)^2=(x-1)^2\) taking sqrt gives \(3x-1=\pm(x-1)\) 3x-1=x-1 or 3x-1=-(x-1)=-x+1 nw try to solve these two equations seperately for x. Does that help?

OpenStudy (anonymous):

yes now I got it I forgot that there are two equations. thanks alot :)

OpenStudy (ajprincess):

Welcome:)